OpenStudy (anonymous):

Which ordered pair is a solution to the equation 3x + 5y = 9? A. (–2, –3) B. (–2, 3) C. (2, –3) D. (2, 3)

OpenStudy (nurali):

3x + 5y = 9 (–2, 3) 3(-2)+5(3)=9 -6+15=9 9=9

OpenStudy (anonymous):

Can you explain please?

OpenStudy (anonymous):

Thanks though.

OpenStudy (anonymous):

Nvm I get it.

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

you are supposed to replace the x and y values with the coordinates.

OpenStudy (nurali):

Anytime

OpenStudy (anonymous):

Oh ok. Thx